OpenStudy (anonymous):

How many solutions are there for the system shown below? x^2 + 4y^2 = 100 4y - x^2 = - 20 A. 4 B. 3 <-- I think it has 3...? C. 1 D. 2

OpenStudy (anonymous):

can u plz show your work?

OpenStudy (anonymous):

4y - x^2 = -20 4 * (-5) - x^2 = -20 -20 - x^2 = -20 0 = x^2 x = 0 I had help for this one ... i had a hard time with it 4y^2 + x^2 + 4y - x^2 = 100 - 20 4y^2 + 4y = 80 y^2 + y = 20 y^2 + y - 20 = 0 y = (-1 +/- sqrt(1 + 80)) / 2 y = (-1 +/- 9) / 2 y = -10/2 , 8/2 y = -5 , 4 4 * 4 - x^2 = -20 16 - x^2 = -20 36 = x^2 x = -6 , 6

OpenStudy (anonymous):

It has three

OpenStudy (anonymous):

Quite Right...:)

OpenStudy (anonymous):

sweet thanks!!!

OpenStudy (anonymous):

yw........:)